OpenStudy (anonymous):

THE SIDES OF THE RIGHT TRIANGLE IS 12 AND 15cm .FIND THE HYPOTENUSE? (with medal)

OpenStudy (nikato):

do you know the pythagorean theorem? a^2 +b^2 =c^2

OpenStudy (anonymous):

\[\sqrt{12^2+15^2}=3 \sqrt{41}=19.2094 \]

OpenStudy (yttrium):

^Such that a and b represents the legs of the rt triangle and c represents the hypotenuse of that triangle.

OpenStudy (yttrium):

And follow what @robtobey did :))